OpenStudy (anonymous):

Hi all, I have one from a sample exam paper I don't know how to approach. I feel stupid for not knowing but I have tried a few things but am getting nowhere. Here it is: "The mass (m) hanging on an elastic band moves according to the equation [m*(d2s/dt2) = - mg/p(s-L)], where (s) is the displacement of the mass, (L) is the natural length of the band and (p) is the constant parameter (called reference elongation). Find s as a function of time subject to the initial conditions: s = s(subnaught), ds/dt = 0 when t = 0. Any help would be vastly appreciated. Regards, Rick

zepdrix (zepdrix):

Is the (s-L) in the denominator? It's a little hard to read it properly with how it's written ^^

OpenStudy (anonymous):

I have a ss of 60 can I help someone with 99 ss

zepdrix (zepdrix):

\[\Large m\cdot s''\quad=\quad -\frac{mg}{p}(s-L)\] Is this what it's suppose to look like Rick? Sorry for my ignorance, I'm not very familiar with physics and all that fancy stuff.

OpenStudy (anonymous):

Hi Zepdrix the equation you did is exactly the one! The (s-L) is not in the denominator.

OpenStudy (anonymous):

Hi "thissucks" I appreciate help from anyone.

zepdrix (zepdrix):

Distributing the -mg/p to the terms in the brackets gives us, \[\Large m s''\quad=\quad -\frac{mg}{p}s+\frac{mgL}{p}\]Adding (mg/p)s to each side gives us,\[\Large m s''+\frac{mg}{p}s\quad=\quad \frac{mgL}{p}\]

zepdrix (zepdrix):

Then I guess we would want to put it into standard form (taking the coefficient off of the leading term) so we can ... hmm that would eliminate all of the m's from the equation though... hmm

zepdrix (zepdrix):

Either that's telling us that the displacement does not depend on the mass of the object, or I made a mistake somewhere lol :) Ok let's just continue, I feel like maybe we're on the right track .

OpenStudy (anonymous):

This is why I find the question hard as I am not exactly sure what it is asking for, that and I am still an amateur at DE's. It says to find s as a function of time under the initial conditions. if that helps?

zepdrix (zepdrix):

The initial conditions will help us `AFTER` we find our general solution. So let's try to achieve that first. Dividing through by m gives us,\[\Large s''+\frac{g}{p}s\quad=\quad \frac{gL}{p}\]We'll have to find the homogeneous solution and then the particular after that.

zepdrix (zepdrix):

So if we look at the homogeneous case:\[\Large s''+\frac{g}{p}s\quad=\quad 0\]

zepdrix (zepdrix):

Are you familiar with solving DE's like this one? Or at least familiar with what the characteristic equation will look like? :) In this case our characteristic equation will be like this:\[\Large r^2+\frac{g}{p}=0\]

OpenStudy (anonymous):

Ah excellent! I kind of now what is going on know, is this a second order non homogenous de?

OpenStudy (anonymous):

I am somewhat familiar. Very new to them.

zepdrix (zepdrix):

Ah ok good :) familiar ground!

zepdrix (zepdrix):

Yes, it's a Second-Order Linear Non-Homogeneous DE

OpenStudy (anonymous):

I know we need to find the overall general solution and that is achieved by setting something to zero and ignoring the rhs, and then finding the particular solution and adding? I am pretty bad at them currently. Which I guess is why I didnt even realise the question was one.

zepdrix (zepdrix):

Well the problem involves a ton of constants, very easy to get lost in it.

OpenStudy (anonymous):

Is this something you are able to do? If not I completely understand and am already in your debt for helping thus far.

zepdrix (zepdrix):

Ya I think we'll be able to get through this one without too much trouble, we'll see how it goes.

zepdrix (zepdrix):

So from the characteristic equation:\[\Large r^2+\frac{g}{p}=0\] Solving for r gives us:\[\Large r\quad=\quad \pm\sqrt{-\frac{g}{p}}\]

zepdrix (zepdrix):

Do you recall which type of solution will give us complex roots in the characteristic equation? Exponentials? Sines/Cosines? Logs? Anything sound familiar? :)

OpenStudy (anonymous):

would it be sine/cosines?

OpenStudy (anonymous):

I know with exponentials we have to sub the r values into each, well i think so anyway.

OpenStudy (anonymous):

I am going to start writing this down so I can refer back to it as this will be a great example, as the couple I did were not as difficult as this.

zepdrix (zepdrix):

yes good good. I guess it technically would be exponentials + sines/cosines. But the exponential portion comes from the `real` part of our r's. And you can see that our r's contain no real, just imaginary parts. So our solution to the homogeneous will contain only sines/cosines.

OpenStudy (anonymous):

Ah I see, this is the part were we are trying to guess what will fit our de the best? Am I on the right track?

zepdrix (zepdrix):

When: \[\Large r=\alpha \pm \beta i\] Our solution is of the form:\[\Large y_h\quad=\quad c_1 e^{\alpha}\cos \beta+c_2 e^{\alpha}\sin \beta\] ^ Just something to remember.

zepdrix (zepdrix):

Fit our de the best? +_+ Hmm

zepdrix (zepdrix):

Don't feel like you have to hurry and write all this down now. The thread will be saved in your profile under the `Questions Asked` section. So you'll always be able to refer back to it.

OpenStudy (anonymous):

Oh excellent, I did not know that.

OpenStudy (anonymous):

Yeah fit the de, our educated guess, well that's how someone described it to me. Wasn't a very great explanation clearly haha

zepdrix (zepdrix):

lol XD

zepdrix (zepdrix):

The roots of our characteristic equation gave us: \[\Large r\quad=\quad 0\pm\sqrt{\frac{g}{p}}i\] So our homogeneous solution will be of the form:\[\Large y_h\quad=\quad c_1 \cos \left(\sqrt{\frac{g}{p}}\;t\right)+c_2 \sin\left(\sqrt{\frac{g}{p}}\;t\right)\]

zepdrix (zepdrix):

And I'm trying to remember ... ummm.. I think we have to find the particular solution BEFORE we can use our initial data. Otherwise we'll run into trouble.

OpenStudy (anonymous):

That makes sense as we will then have all the required forms and then we can just sub in the IC?

zepdrix (zepdrix):

Yes, and hopefully the IC will allow us to solve for the missing constants, c_1 and c_2, giving us a `specific` solution instead of the general form.

OpenStudy (anonymous):

Awesome

zepdrix (zepdrix):

\[\Large s''+\frac{g}{p}s\quad=\quad \frac{gL}{p}\] Our particular solution will be of the same form that the `right side` of the equation has. Maybe this is what you were referring to by "guessing what will fit our DE best".

OpenStudy (anonymous):

yes I think that is! I was shown that we need to match the constants up and then solve it, but I don't remember much else.

zepdrix (zepdrix):

Since our `right hand side` is simply constant, the form of the particular solution will also be constant.\[\Large y_p\quad=\quad C\]\[\Large y_p'\quad=\quad 0\]\[\Large y_p''\quad=\quad 0\] So I set the y_p (particular solution) as an arbitrary constant and then took it's derivative a couple times. We want to use this list of information and plug it back into the original equation to solve for our constant C.

zepdrix (zepdrix):

Oh I'm calling these y's... my bad. I hope that's not too confusing.

zepdrix (zepdrix):

I tend to mix up the letters sometimes lol.

OpenStudy (anonymous):

no thats all good, is that convention as that is what I have seen? :)

zepdrix (zepdrix):

I'm just so used to seeing these types of equations as y''+y'+... instead of involving s. This is kind of a neat problem though :x cool physics stuff. So here is our homogeneous part of the solution: \[\Large s_h\quad=\quad c_1 \cos \left(\sqrt{\frac{g}{p}}\;t\right)+c_2 \sin\left(\sqrt{\frac{g}{p}}\;t\right)\] I just wanted to fix the y a sec there.

OpenStudy (anonymous):

I am glad you are enjoying it :D I am enjoying learning!

OpenStudy (anonymous):

This will help me heaps in a couple of weeks when I have my first exam

zepdrix (zepdrix):

\[\Large s_p\quad=\quad C\]\[\Large s_p''\quad=\quad 0\] Plugging these into our original equation gives us:\[\Large s_p''+\frac{g}{p}s_p\quad=\quad \frac{gL}{p}\qquad\quad\to\quad\qquad 0+\frac{g}{p}C\quad=\quad \frac{gL}{p}\]

zepdrix (zepdrix):

So what is our particular solution?? :) It should be jumping out at you! hehe

OpenStudy (anonymous):

Is it L? :/ I feel like an idiot. Hold on give me a sec

OpenStudy (anonymous):

C = L?

zepdrix (zepdrix):

Yes, good good! C=L And s_p=C is our particular solution.

OpenStudy (anonymous):

No that isn't it

OpenStudy (anonymous):

oh sweet!

zepdrix (zepdrix):

\[\Large s_h\quad=\quad c_1 \cos \left(\sqrt{\frac{g}{p}}\;t\right)+c_2 \sin\left(\sqrt{\frac{g}{p}}\;t\right)\]\[\Large s_h\quad=\quad L\] And as you mentioned before, our general solution will be the sum of these, yes?\[\Large s\quad=\quad s_h+s_p\]

zepdrix (zepdrix):

\[\Large s_p\quad=\quad L\]Grr typo..

OpenStudy (anonymous):

Yeah oh awesome, so the general solution will be :

OpenStudy (anonymous):

c1cos(gp−−−√t)+c2sin(gp−−−√t) + L

OpenStudy (anonymous):

Wait that didn't come out right haha

zepdrix (zepdrix):

ah poor guy XD

OpenStudy (anonymous):

haha so my answer is then the above which is S(t) = s_p +s_h

zepdrix (zepdrix):

brb need some chocolate milk from the kitch:x

OpenStudy (anonymous):

That is awesome, that is a very long drawn out problem. So in a physics sense it is telling me then that the displacement as a function of time is governed by the length, obviously, and the constant parameter.

OpenStudy (anonymous):

haha okay, get me a glass while you're there ;D

zepdrix (zepdrix):

Ok good so the sum of our parts gives us:\[\Large s(t)\quad=\quad c_1 \cos \left(\sqrt{\frac{g}{p}}\;t\right)+c_2 \sin\left(\sqrt{\frac{g}{p}}\;t\right)+L\] Something like that, right?

zepdrix (zepdrix):

And our initial conditions are:\[\Large s(0)\quad=\quad s_o\]\[\Large s'(0)\quad=\quad0\]

OpenStudy (anonymous):

Yep that is them, so we don't need to do anything else with it do we? As with s(0) = s_0 would just equal L?

zepdrix (zepdrix):

Hmm let's see.. I think it gives us a little bit more than that.

OpenStudy (anonymous):

oh wait one sec

OpenStudy (anonymous):

if t = 0 then it would be ccos(0) + csin(0) + L

zepdrix (zepdrix):

Hmm cos(0)=?

OpenStudy (anonymous):

Wait... that was stupid sorry

OpenStudy (anonymous):

C_1 + C_2 + L as cos (0) is 1? I am most probably wrong

OpenStudy (anonymous):

My thinking was that due to t = 0 that would cancel out the fraction inside the cos (*)

zepdrix (zepdrix):

Ya it kills the fraction inside of cosine and sine, making 0's inside those spots. cos(0)=1, sin(0)=0. So it looks like s(0) is telling us something about c_1, right?\[\Large s_o\quad=\quad c_1 \cos(0)+\cancel{c_2\sin(0)}+L\]

OpenStudy (anonymous):

oh damn sorry I completely forgot the sin!

OpenStudy (anonymous):

so if we rearrange for C_1

OpenStudy (anonymous):

it is equal to 1? or s(t) - L

zepdrix (zepdrix):

Ya you're on the right track, your notation is a little sloppy though ^^ \[\Large s(0)\quad=\quad s_o\quad = \quad c_1+L\]So solving for c_1 gives us:\[\Large c_1\quad=\quad s_o-L\]

OpenStudy (anonymous):

Yeah it is terrible, I haven't really done the whole typing equations out a lot, I am a pen and paper guy haha

zepdrix (zepdrix):

Ya getting math help online can be a real pain lol. I LOVE this equation tool now that I've finally gotten used to it. Takes some time though.

OpenStudy (anonymous):

I will definitely have to keep using it as the help you have given today is better than I got at uni from a guy that gets paid, you should do tutoring if you already are not.

OpenStudy (anonymous):

So can we do anything else with it? We now have the constant, there is not much more we could do is there? Or maybe we need to take the derivative? I don't know

zepdrix (zepdrix):

Ya I have an interview new week actually :) Trying to find some tutoring work while I'm in school. This math is all fresh in my mind so it's a good time to try.

zepdrix (zepdrix):

Yes. So let's see if the other initial condition tells us anything about c_2.

zepdrix (zepdrix):

Derivative time.

zepdrix (zepdrix):

\[\Large s(t)\quad=\quad c_1 \cos \left(\sqrt{\frac{g}{p}}\;t\right)+c_2 \sin\left(\sqrt{\frac{g}{p}}\;t\right)+L\] \[\Large s'(t)\quad=\quad ?\]

zepdrix (zepdrix):

Don't forget your chain rule! This one won't be too bad.

OpenStudy (anonymous):

be one sec :)

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

Did it go through?

zepdrix (zepdrix):

If it's too much of a pain to type out, that's fine. I can do it lol

zepdrix (zepdrix):

It didn't :OOO

OpenStudy (anonymous):

nah ill type it out again only be a sec

OpenStudy (anonymous):

\[\sqrt{g/p}(\cos(t \sqrt{g/p})) - \sin(t \sqrt{g/p})\]

OpenStudy (anonymous):

I think thats it.

OpenStudy (anonymous):

I know i should know derivatives of cos and sin and the chain rule inside out, but I am a bit rusty.

zepdrix (zepdrix):

Ok looks pretty close:\[\Large s'(t)\quad=\quad \color{red}{c_2}\sqrt{g/p}(\cos(t \sqrt{g/p})) -\color{red}{c_1}\color{red}{\sqrt{g/p}}\sin(t \sqrt{g/p})\]